OpenStudy (anonymous):

let f(x)=3x-10 find the value of f(-3)

OpenStudy (anonymous):

f(-3)=3(-3)-10 =-9-10 =-19 f(-3) = -19

OpenStudy (anonymous):

Does that answer your question?